# Determinant Estimation under Memory Constraints and Neural Scaling Laws

### v1

@@ 3.1 Neural Tangent Kernels and Scaling Laws
Let $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a positive-definite kernel. For a sequence of inputs $\{x_{i}\}_{i=1}^{\infty}\subset\mathcal{X}$ and for each $n=1,2,\dots$, let $\mathbf{K}_{n}\coloneqq(\kappa(x_{i},x_{j}))_{i,j=1}^{n}$ be the corresponding Gram matrix for the first $n$ inputs. Our objective is to estimate $\mathsf{logdet}(\mathbf{K}_{n})$ for large $n$, using computations involving only smaller values of $n$. Recall that $f$ is a Gaussian process with mean function $\mu:\mathcal{X}\to\mathbb{R}^{d}$ and covariance kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, denoted $f\sim\mathcal{GP}(\mu,\kappa)$, if for every $x_{1},\dots,x_{n}\in\mathcal{X}$ and $d\geq 1$, $(f(x_{i}))_{i=1}^{n}\sim\mathcal{N}((\mu(x_{i}))_{i=1}^{n},(\kappa(x_{i},x_{j}))_{i,j=1}^{n})$. Our results are founded on the following lemma, which we prove in Appendix G.
@@
where $r$ is the rank of $\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1)$. In the case where $\mathrm{rank}(\mathbf{K}_{n})-\mathrm{rank}(\mathbf{K}_{n-1})=d$ this reduces to
@@
Lemma 1 is particularly interesting since it highlights a connection between the determinants of Gram matrices and error curves for GPs. This bounds the effect of adding or removing a datapoint on the determinant in terms of the prior variance of a corresponding GP. Previous theoretical (see Appendix F) and empirical studies (Spigler et al., , 2020; Bahri et al., , 2021; Li et al., , 2023; Barzilai & Shamir, , 2023) have established a power law relationship of the form $E(n)=\Theta(n^{-\xi})$ as $n\to\infty$ for some $\xi>0$. Invoking Lemma 1 in view of this, we propose the following scaling law for determinants, demonstrated in Figure 2.
@@
Under Assumption 1 and Assumption 2, we derive an expression for the asymptotic behavior of the normalized log-determinants. The proof is given in Appendix G.
@@
  • (LLN) Under Assumption 1, there exist constants $c_{0},\nu>0$ such that as $n\to\infty$,
@@ 3.2 FLODANCE
As (4) is linear in the unknown parameters $c_{0},\nu$, these parameters can be estimated by linear regression on a precomputed sequence  [^2] This is generally a more stable regression problem than estimating $C$ and $\nu$ directly from (3). $(L_{1},\dots,L_{n})$. This sequence can be computed using a single pass of MEMDET on a subsample of the Gram matrix. After performing linear regression, by extrapolating to larger $n$, the normalized log-determinants of larger NTK Gram matrices can be estimated. The ordering of the data is arbitrary, but will affect the output of the regression task. We refer to this as the Factorial-based Log-Determinant Analysis and Numerical Curve Estimation procedure, or FLODANCE. To improve performance, we allow for a non-asymptotic correction to the exponent $\nu$ as $\nu(n)=\nu_{0}+\sum_{i=1}^{q}\nu_{i}n^{-i}$ (in practice, we find $q=3\sim 6$ to work well). In light of (5) and discussion in Appendix H, for large $n$, we expect that approximately
$y_{n}=c_{0}x_{n,0}+\sum_{i=1}^{q+1}\nu_{i-1}x_{n,i}+\epsilon_{n},\quad\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2}),$
where $y_{n}=\frac{n}{\sqrt{n-1}}(L_{n}-L_{1})$, $x_{n,0}=\sqrt{n-1}$, and $x_{n,i}=\frac{-\log(n!)}{n^{i-1}\sqrt{n-1}}$ for $i=1,...,m+1$ are the covariates. The corresponding numerical procedure is presented in Algorithm 1 and the implementation can be found in LABEL:list:detkit2.
@@
Number of model’s output dimension $d$ (i.e., classes or labels)
Total number of datapoints $n$ (i.e., training set size),
Number of datapoint samples $n_{s}$ where $1<n_{s}\leq n$,
Number of burn-in terms $n_{0}$ where $1\leq n_{0}<n_{s}$,
@@
Output : Estimated normalized log-determinant $\hat{L}_{n}$ of matrix $\mathbf{K}_{n}$ of size $m\times m$ where $m=nd$
@@
// Denote $\mathbf{K}_{n_{s}[:k,:k]}$ the $k\times k$ principal sub-matrix of $\mathbf{K}_{n_{s}}$
@@
// Normalize log-determinants and record every $d$-th entry in the sequence.
5 Obtain $(L_{j})_{j=n_{0}}^{n_{s}}$ where $L_{j}\leftarrow j^{-1}\ell_{m_{j}}$ and $m_{j}\leftarrow jd$.
@@ 11      $X_{j,1}\leftarrow(n_{j}-1)^{\frac{1}{2}}$
              // $\operatorname{log\Gamma}(n+1)=\log(n!)$ is the Log-gamma function.
@@
// Perform linear regression $\boldsymbol{y}=\mathbf{X}\boldsymbol{\beta}+\epsilon$ to estimate coefficients $\boldsymbol{\beta}$
@@
In practice, we also observe that a burn-in period may be required to obtain accurate estimates of $c_{0}$ and the $\nu_{i}$ that appear in Proposition 1. Better performance was often achieved in our experiments by discarding the early determinant samples, effectively replacing the $L_{1}$ term appearing in $y_{n}$ with different constants $L_{k+1}$, for a burn-in of length $k$. This appears to be due to the sudden appearance of very small eigenvalues that shift the model fit, and appears to be a consistent phenomena that warrants further investigation. We found that when needed, the burn-in required was always less than 500 terms, verified by cross-validation.
Table 2: Comparison of various approximations of the normalized log-determinant $\hat{L}_{n}$ with the exact computation $L_{n}$ obtained in 64-bit floating-point precision (first row). Values represent averages over five trained networks, with standard deviations in parentheses. Bold values indicate the closest approximation, with the next-best underlined.
Quantity Model Configuration ResNet9 ResNet9 ResNet18 MobileNet Dataset CIFAR-10 CIFAR-10 CIFAR-10 MNIST Subsample Size $n=1000$ $n=2500$ $n=1000$ $n=2500$ $L_{n}$ Direct Computation (64-bit) (Reference) $-0.490$ ($0.20$) $-4.48$ ($0.26$) $-15.5$ ($0.84$) $-81.0$ ($3.15$) Approximation $(\hat{L}_{n})$ Direct Computation (16-bit) $9.00$ ($0.16$) $7.89$ ($0.21$) $-6.32$ ($1.07$) $-26.7$ ($0.85$) Direct Computation (32-bit) $2.32$ ($0.18$) $0.43$ ($0.22$) $-12.1$ ($0.83$) $\boldsymbol{-69.4}$ ($2.00$) Block Diagonal $18.0$ ($0.13$) $18.0$ ($0.17$) $4.77$ ($1.06$) $10.2$ ($1.40$) SLQ $14.2$ ($0.15$) $13.9$ ($0.22$) $-0.66$ ($1.06$) $-12.5$ ($1.37$) Pseudo NTK $-13.9$ ($0.05$) $-16.2$ ($0.15$) $\boldsymbol{-14.89}$ ($0.15$) $-19.9$ ($0.19$) FLODANCE $n_{0}=1,n_{s}=50$ $4.32$ ($0.46$) $1.96$ ($0.69$) $-8.40$ ($1.20$) $-55.0$ ($2.20$) FLODANCE $n_{0}=1,n_{s}=100$ $3.31$ ($0.27$) $0.91$ ($0.54$) $-10.6$ ($0.70$) $-54.0$ ($2.06$) FLODANCE $n_{0}=300,n_{s}=500$ $\boldsymbol{0.53}$ ($0.27$) $\boldsymbol{-3.50}$ ($0.33$) $-13.6$ ($0.43$) $-62.3$ ($1.13$)
@@ 4 Numerical Experiments
We are now in a position to test the accuracy of the FLODANCE algorithm for estimation of log-determinants on large-scale problems of interest. The test problems that we consider are NTK matrices corresponding to common deep learning models: ResNet9, ResNet18, and ResNet50 (He et al., , 2016) trained on the CIFAR-10 dataset (Krizhevsky, , 2009), and MobileNet (Howard et al., , 2017) trained on the MNIST dataset (LeCun et al., , 1998). Our experiments are split into two sections: first, the dataset size is reduced in order to enable the matrices to fit into memory on a consumer device, in Section 4.1; and then larger subsamples and full datasets are considered, in Section 4.2.333Experiments in this section were conducted on a desktop-class device with an AMD Ryzen®7 5800X processor, NVIDIA RTX 3080, and 64GB RAM. As a baseline, we employ MEMDET to compute the relevant quantities, where the accuracy is limited only by numerical precision.
@@ 4.1 Smaller Data Sets
In order to demonstrate the scaling laws for ratios of successive determinants of NTK Gram matrices, a ResNet50 model was trained on a subset of 1000 datapoints from the CIFAR-10 dataset. The resulting matrix $\mathbf{K}_{1000}$ is of size $10000$\times$10000$$. In Figure 2, we plot exact ratios $\det(\mathbf{K}_{n})/\det(\mathbf{K}_{n-1})$, with the blue line representing the line of best fit under the scaling law Assumption 1.
For baseline comparison to existing techniques, we compared FLODANCE to approximations of both the matrices themselves and their log-determinants. For the matrix approximations, we consider a block-diagonal approximation ignoring between-data correlations, as well as the pseudo-NTK matrix studied in (Mohamadi et al., , 2023). As an approximate log-determinant technique, we consider the stochastic Lanczos quadrature, often regarded as the current state-of-the-art in log-determinant estimation of large matrices (Gardner et al., , 2018). We further compare the exact methods across 16-, 32- and 64-bit (treated as exact) floating-point precision, in order to compare the accuracy of our extrapolation method against memory-saving mixed-precision calculations.
To this end, ResNet9 was trained on a subsets of 1000 and 2500 images from CIFAR-10, and ResNet18 on 1000 datapoints. MobileNet was also trained on a 2500 image subset of the MNIST dataset. A comparison of the different methods is presented in Table 2. We see that all existing NTK and log-determinant approximation techniques perform poorly when compared with 16- and 32-bit mixed-precision calculations. On the other hand, the FLODANCE estimates that contained a burn-in phase consistently either outperformed the mixed-precision approximations or were competitive. When no burn-in was used, FLODANCE still outperformed the approximation methods. This suggests that at this scale, the error in the scaling law approximation is less than the mixed-precision errors discussed in Section 2.1. Extrapolating determinants based on expected behavior can bypass numerical issues at scale. Comparisons for ResNet9 trained on 2500 images from the CIFAR-10 dataset are visualized in Figure 4. FLODANCE estimates consistently outperform the competing methods.
@@
Figure 5: Log-determinant $\ell_{n}$ for $n=1,\dots,$50000$$, corresponding to $m\times m$ NTK submatrices where $m=nd$ and $d=10$, from ResNet9 (a) and ResNet50 (b) trained on CIFAR-10 with $50000$ datapoints. Values are computed using MEMDET (Algorithm 3) with LDL decomposition (black curves, overlaid by colored curves). FLODANCE (Algorithm 1) is fitted in a small region (shaded gray) and extrapolated over a much larger interval. The yellow curve in the interval $(n_{0},n_{s})=(10^{2},5\times 10^{3})$ represents the fit, while the red curve in $(n_{s},n)=(5\times 10^{3},5\times 10^{4})$ shows the extrapolation. The blue curves, corresponding to the right axis in each panel shows the relative error, reaching impressive 0.05% in (a) and 0.02% in (b).

### v2

@@ 3.1 Neural Tangent Kernels and Scaling Laws
Let $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}^{d\times d}$ be a positive-definite kernel. For a sequence of inputs $\{x_{i}\}_{i=1}^{\infty}\subset\mathcal{X}$ and for each $n\in\mathbb{N}$, let $\mathbf{K}_{n}\coloneqq(\kappa(x_{i},x_{j}))_{i,j=1}^{n}$ be the corresponding Gram matrix for the first $n$ inputs, which is a matrix of size $nd\times nd$. Our objective is to estimate $\mathsf{logdet}(\mathbf{K}_{n})$ for large $n$, using computations involving only smaller values of $n$. Recall that $f$ is a Gaussian process with mean function $\mu:\mathcal{X}\to\mathbb{R}^{d}$ and covariance kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathbb{R}^{d\times d}$, denoted $f\sim\mathcal{GP}(\mu,\kappa)$, if for every $x_{1},\dots,x_{n}\in\mathcal{X}$ and $d\geq 1$, $(f(x_{i}))_{i=1}^{n}\sim\mathcal{N}((\mu(x_{i}))_{i=1}^{n},(\kappa(x_{i},x_{j}))_{i,j=1}^{n})$. Our results are founded on the following lemma, which we prove in Section F.2.
@@
where $r$ is the rank of $\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1)$. In the case where $\mathrm{rank}(\mathbf{K}_{n})-\mathrm{rank}(\mathbf{K}_{n-1})=d$, this reduces to
@@
Lemma 1 is particularly interesting since it highlights a connection between the determinants of Gram matrices and error curves for GPs. This bounds the effect of adding or removing a datapoint on the determinant in terms of the prior variance of a corresponding GP. Previous theoretical (see Section F.1) and empirical studies (Spigler et al., , 2020; Bahri et al., , 2024; Li et al., , 2023; Barzilai & Shamir, , 2024) have established a power law relationship of the form $E(n)=\Theta(n^{-\xi})$ as $n\to\infty$ for some $\xi>0$. In view of this, and invoking Lemma 1, we propose the following scaling law for determinants, demonstrated in Figure 2.
@@
Under Assumption 1 and Assumption 2, we derive an expression for the asymptotic behavior of the normalized log-determinants. The proof is given in Section F.2.
@@
  • (LLN) Under Assumption 1, there exist constants $c_{0},~{}\nu~{}>~{}0$ such that as $n\to\infty$,
@@ 3.2 FLODANCE
As (4) is linear in the unknown parameters $c_{0},\nu$, these parameters can be estimated by linear regression on a precomputed sequence $(L_{1},\dots,L_{n})$. [^1] This is generally a more stable regression problem than estimating $C$ and $\nu$ directly from (3). This sequence can be computed using a single pass of MEMDET on a subsample of the Gram matrix. After performing linear regression, by extrapolating to larger $n$, the normalized log-determinants of larger NTK Gram matrices can be estimated. The ordering of the data is arbitrary, but it will affect the output of the regression task. We refer to this method as the Factorial-based Log-Determinant Analysis and Numerical Curve Estimation procedure, or FLODANCE. To improve performance, we allow for a non-asymptotic correction to the exponent $\nu$ as $\nu(n)\coloneqq\nu_{0}+\sum_{i=1}^{q}\nu_{i}n^{-i}$ (in practice, we find $q\leq 10$ works well). In light of (5) and discussion in Section F.3, for large $n$, we expect that approximately
$y_{n}=c_{0}x_{n,0}+\sum_{i=1}^{q+1}\nu_{i-1}x_{n,i}+\epsilon_{n},\quad\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2}),$ (6)
where $y_{n}\coloneqq\frac{n}{\sqrt{n-1}}(L_{n}-L_{1})$, $x_{n,0}\coloneqq\sqrt{n-1}$, and $x_{n,i}\coloneqq\frac{-\log(n!)}{n^{i-1}\sqrt{n-1}}$ for $i=1,...,m+1$ are the covariates. The corresponding numerical procedure is presented in Algorithm 1 and the implementation can be found in LABEL:list:detkit2.
@@
Model’s output dimension $d$,
Total number of datapoints $n$,
Data subsample size $1<n_{s}\leq n$,
Burn-in length $1\leq n_{0}<n_{s}$,
@@
Output : Estimated normalized log-determinant $\hat{L}_{n}$ of $\mathbf{K}_{n}$ of size $m\times m$ where $m=nd$
@@
// $\mathbf{K}_{n_{s}[:k,:k]}$ is the $k\times k$ principal sub-matrix of $\mathbf{K}_{n_{s}}$
@@
// Normalize and record every $d$-th entry.
5 Obtain $(L_{j})_{j=n_{0}}^{n_{s}}$ for $L_{j}\leftarrow j^{-1}\ell_{m_{j}}$ and $m_{j}\leftarrow jd$.
@@ 11      $X_{j,1}\leftarrow(n_{j}-1)^{\frac{1}{2}}$
              // The Log-gamma function computes $\log(n_{j}!)$.
@@
// Estimate regression coefficients $\boldsymbol{\beta}$ in $\boldsymbol{y}=\mathbf{X}\boldsymbol{\beta}+\epsilon$
@@
In practice, we also observe that a burn-in period may be required to obtain accurate estimates of $c_{0}$ and the $\nu_{i}$ that appear in Proposition 1. Better performance was often achieved in our experiments by discarding the early determinant samples, effectively replacing the $L_{1}$ term appearing in $y_{n}$ with different constants $L_{n_{0}}$, for a burn-in of length $n_{0}-1$. This seems to be due to the sudden emergence of very small eigenvalues that shift the model fit, and constitutes a consistent phenomenon that warrants further investigation. We found that when needed, the burn-in required was always less than 500 terms, verified by cross-validation.
@@ 4 Numerical Experiments
We now evaluate the accuracy of the FLODANCE algorithm for estimating log-determinants on large-scale problems of interest. The test problems that we consider are NTK matrices corresponding to common deep learning models: ResNet9, ResNet18, and ResNet50 (He et al., , 2016) trained on the CIFAR-10 dataset (Krizhevsky, , 2009), and MobileNet (Howard et al., , 2017) trained on the MNIST dataset (LeCun et al., , 1998). Our experiments are split into two sections: first, the dataset size is reduced in order to enable the matrices to fit into memory on a consumer device, in Section 4.1; and then larger subsamples and full datasets are considered, in Section 4.2.222Experiments in this section were conducted on a desktop-class device with an AMD Ryzen®7 5800X processor, NVIDIA RTX 3080, and 64GB RAM. As a baseline, we employ MEMDET to compute the relevant quantities, where accuracy is limited only by numerical precision. Detailed runtime and performance diagnostics for MEMDET are provided in Appendix E. Additional experiments evaluating the performance of FLODANCE appear in Appendix G, including its robustness (Section G.1) and its application to Matérn kernel Gram matrices (Section G.2).
@@ 4.1 Smaller Data Sets
In order to demonstrate the scaling laws for ratios of successive determinants of NTK Gram matrices, a ResNet50 model was trained on a subset of 1000 datapoints from the CIFAR-10 dataset with $d=10$ classes. The resulting matrix $\mathbf{K}_{1000}$ is of size $10000$\times$10000$$. In Figure 2, we plot exact ratios $\det(\mathbf{K}_{n})/\det(\mathbf{K}_{n-1})$, with the blue line representing the line of best fit under the scaling law Assumption 1.
For baseline comparison to existing techniques, we compared FLODANCE to approximations of both the matrices themselves and their log-determinants. For the matrix approximations, we consider a block-diagonal approximation ignoring between-data correlations, as well as the pseudo-NTK matrix studied in (Mohamadi et al., , 2023). As an approximate log-determinant technique, we consider stochastic Lanczos quadrature, often regarded as the current state-of-the-art for large-scale log-determinant estimation (Gardner et al., , 2018). We also compare exact methods across 16-, 32-, and 64-bit (treated as exact) floating-point precision, to assess the accuracy of our extrapolation against memory-saving mixed-precision calculations.
To this end, ResNet9 was trained on a subsets of 1000 and 2500 images from CIFAR-10, and ResNet18 on 1000 datapoints. MobileNet was also trained on a 2500 image subset of the MNIST dataset with $d=10$ classes. A comparison of the different methods is presented in Table 2. We see that all existing NTK and log-determinant approximation techniques perform poorly when compared with 16- and 32-bit mixed-precision calculations. On the other hand, the FLODANCE estimates that contained a burn-in phase consistently either outperformed the mixed-precision approximations or were competitive. When no burn-in was used, FLODANCE still outperformed the approximation methods. This suggests that at this scale, the error in the scaling law approximation is less than the mixed-precision errors discussed in Section 2.1. Extrapolating determinants based on expected behavior can bypass numerical issues at scale. Comparisons for ResNet9 trained on 2500 images from the CIFAR-10 dataset are visualized in Figure 4. FLODANCE estimates consistently outperform the competing methods.
@@
Table 2: Comparison of approximations of the log-determinant ^$\hat{\ell}_{n}$ with the exact computation $\ell_{n}$ obtained in 64-bit floating-point precision (first row). Values represent average percentage relative errors over five trained networks, with standard deviations in parentheses. Bold values indicate the closest approximation, with the next-best underlined. For the corresponding compute times, see Table H.2.
Quantity Model $\rightarrow$ Configuration ResNet9 ResNet9 ResNet18 MobileNet Dataset CIFAR-10 CIFAR-10 CIFAR-10 MNIST Subsample Size $n=1000$ $n=2500$ $n=1000$ $n=2500$ $\ell_{n}$ Direct Computation (64-bit) (Reference) $76538$ ($203$) $181377$ ($649$) $65630$ ($842$) $-183962$ ($7869$) Relative Error |^$\frac{|\hat{\ell}_{n}-\ell_{n}|}{\ell_{n}}$ Direct Computation (16-bit) $12.41\%$ ($0.12$) $17.05\%$ ($0.13$) $14.00\%$ ($0.24$) $66.97\%$ ($2.13$) Direct Computation (32-bit) $3.67\%$ ($0.06$) $6.77\%$ ($0.08$) $5.25\%$ ($0.09$) $\boldsymbol{14.27}\%$ ($0.95$) SLQ $81.51\%$ ($0.16$) $80.89\%$ ($0.24$) $101.03\%$ ($1.64$) $84.52\%$ ($1.51$) Block Diagonal $76.49\%$ ($0.12$) $75.15\%$ ($0.16$) $92.76\%$ ($1.55$) $112.55\%$ ($1.22$) Pseudo NTK $118.35\%$ ($0.10$) $122.35\%$ ($0.27$) $122.95\%$ ($0.25$) $75.32\%$ ($1.04$) FLODANCE $n_{0}=1$,00 $n_{s}=50$ $7.75\%$ ($0.77$) $11.27\%$ ($1.10$) $12.19\%$ ($0.30$) $36.41\%$ ($2.53$) FLODANCE $n_{0}=1$,00 $n_{s}=100$ $5.61\%$ ($0.32$) $8.54\%$ ($0.63$) $8.09\%$ ($0.68$) $35.51\%$ ($1.46$) FLODANCE $n_{0}=300$, $n_{s}=500$ $\boldsymbol{1.34}\%$ ($0.11$) $\boldsymbol{1.37}\%$ ($0.14$) $\boldsymbol{2.9}\%$ ($0.81$) $23.19\%$ ($1.76$)
Figure 5: Log-determinant $\ell_{n}$ for $n=1,\dots,$50000$$, corresponding to $m\times m$ NTK submatrices where $m=nd$ and $d=10$, from 64-bit NTK matrices of ResNet9 (a) and ResNet50 (b) trained on CIFAR-10 with $50000$ datapoints. Values are computed using MEMDET (Algorithm 3) with LDL decomposition (black curves, overlaid by colored curves). FLODANCE (Algorithm 1) is fitted in a small region (shaded gray) and extrapolated over a much larger interval. The yellow curve in the interval $(n_{0},n_{s})=(10^{2},5\times 10^{3})$ represents the fit, while the red curve in $(n_{s},n)=(5\times 10^{3},5\times 10^{4})$ shows the extrapolation. The blue curves, corresponding to the right axis in each panel shows the relative error, reaching impressive 0.2% in (a) and 0.02% in (b).